OpenStudy (anonymous):

How to simplify (x-9x(x^1/2)) / (x^1/2)

OpenStudy (anonymous):

\[x-9x \sqrt{x}/\sqrt{x}\]

OpenStudy (anonymous):

is this \[\frac{ x - 9 x ^{3/2} }{ x ^{1/2} }\] ??

OpenStudy (anonymous):

ok then how would you take the derivative of that ?

OpenStudy (anonymous):

see x can be written as \[x = (\sqrt{x})^{2}\] now take root of x common from numerator terms and cancel with denominator

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

can someone still show me how to do the derivative of it?

OpenStudy (anonymous):

then the expression will become \[\sqrt{x} - 9 x\]

OpenStudy (anonymous):

\[\sqrt{x}-9\sqrt{x} \right?\]

OpenStudy (anonymous):

\[\sqrt{x}-9\sqrt{x}\] **

OpenStudy (anonymous):

differentiation of x wrt x is 1

OpenStudy (anonymous):

oooh wow ok i got it now!!

OpenStudy (anonymous):

thanks !! just didnt see it lol..

OpenStudy (anonymous):

\[\frac{ d(\sqrt{x}) }{ dx } = \frac{ 1 }{ 2\sqrt{x} }\]